OpenStudy (anonymous):

√(x+2) -2=x solve for x

OpenStudy (nurali):

√(x+2) -2=x √(x+2) =x+2 squaring both side (x+2)=(x+2)^2 x+2=x^2+4x+4 x^2+4x+4-x-2=0 x^2+3x+2=0

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

wouldnt (x+2)^2 get x^2+4

OpenStudy (anonymous):

oh wait im stupid your correct

OpenStudy (anonymous):

so x= -1 and -2

OpenStudy (nurali):

Yes

OpenStudy (anonymous):

Thanks!!

OpenStudy (nurali):

Anytime.